OpenStudy (anonymous):

Which ordered pairs from the replacement set are solutions to the equation? y = -2x + 9 ; Replacement Set {((-1, 11), (0, 8), (1, 7), (2, 5)} Answer (-1, 11) (0, 8) (1, 7) (2, 5)

OpenStudy (mathstudent55):

Take an ordered pair; let's start with the first one: (-1,11) Plug the x-value of the ordered pair (-1) into the x of the equation: y = -2x + 9 y = -2(-1) + 9 =2 + 9 = 11 You get y = 11. Since this ordered pair has 11 as its y value, it is a solution to the equation. Now try the other ordered pairs the same way.